# GlucOS: Security, correctness, and simplicity for automated insulin delivery

### v1

@@ 4.2 Reactive safe model
One wrinkle we add to this otherwise standard calculation is that we learn the physiological parameters we use in these calculations using a data-driven approach for safety. Our learning system consists of a weekly learning task that evaluates three weeks worth of data to infer broader physiological trends. These learned settings become the baseline for our parameters. Then, we use a PID controller from feedback control [22] to refine these values given recent observations. Being able to adapt is important for handling pump malfunction and pump insertion site safety violations (Section 3.2).
Calculating insulin is a simple calculation to implement, consisting of only 25 lines of formally verified code in the BioKernel (Section 8). However, despite this simplicity, it is effective (Sections 10 and 11). The main drawback of this model is that is it unable to detect underdosing insulin during food digestion quickly enough to avoid high glucose.
4.3 Predictive model
To predict insulin underdosing during digestion, we use a food prediction deep-learning model from the literature [45]. Our version of this model is a generalization where we model any source of glucose, not just food, but the details of it are beyond the scope of this paper. In our evaluation our model outperforms the reactive safe model both in simulation (Section 10) and for the individual that uses our system for their real-world insulin dosing (Section 11), showing why it is important for automated insulin delivery systems to have a practical way to adopt ML.
Figure 1 shows an example of when the predictive ML model correctly detects digestion and doses insulin to maintain balance. This figure shows a screenshot from the BioKernel’s main user interface. The individual ate breakfast at 9 am and their glucose levels rose. The predictive ML model correctly detects that the individual is going to absorb additional glucose from digestion, resulting in a predicted glucose level rise over the next hour. Accordingly, the predictive ML model injects enough insulin to cover the rising glucose from the meal and keep the individual’s glucose levels in range.
@@ 5 Driver security: Withstanding malicious pump drivers
Pump drivers provide abstractions for insulin delivery applications to communicate with insulin pumps. Much like traditional device drivers, they have complex implementations since they help accomplish a wide range of tasks including pump setup/teardown, managing pump UX, handling insulin delivery/suspension, managing pump errors, accurately tracking insulin doses etc. Different pump drivers use different implementations to carry out these tasks.
While CVEs do not currently exist for pump drivers, there is strong possibility for bugs or vulnerabilities in their implementation given their complexity and the fact that they not been formally verified. Since drivers control insulin pumps, these vulnerabilities can even lead to the death of individuals. For example, consider a scenario where the user decides to inject 3 U of insulin using the BioKernel. The BioKernel would relay this as a command to the pump through abstractions provided by the pump driver. If the pump driver inadvertently or intentionally issues this command twice, 6 U of insulin would be delivered to the user, which would put the user in danger by drastically reducing the amount of glucose in their body.
With GlucOS, our goal is to support their complexity by allowing users to employ pump drivers of their choice while upholding security. Our security goals are to not just protect users from inadvertent bugs in pump drivers, but to also protect them from malicious pump drivers which could intentionally add, modify or drop pump commands.
@@
5.1 Reference monitor overview
We introduce a reference monitor as part of the BioKernel to handle untrusted pump drivers, which interposes on communications between the pump driver and the pump. We choose to vet these communications instead of implementing our own pump driver to give users the ability to use any pump driver of their choice. This also helps keep our trusted computing base small thereby easing formal verification.
Figure 4 provides an overview of our reference monitor. We only validate commands pertaining to insulin delivery since they pose the most dangerous threats to the user. From a high-level, the pump registers insulin commands with the reference monitor before issuing them to the pump driver. In the reference monitor, we check to ensure that the command issued by the pump driver is consistent with the command registered by the BioKernel. At the reference monitor, we also intercept all records of delivered insulin doses to ensure their integrity for closed loop calculations. On detecting manipulation from the pump driver, we transition insulin delivery to a default safe state (described in Section 7) where there is no immediate threat to the user. We do not validate communication pertaining to pump setup / UX since users can directly detect the respective manipulations. Figure 5 lists the different types of messages pertaining to insulin delivery exchanged between a pump driver and an insulin pump. We list other message types in Appendix B.
5.2 Design Principles
Avoid programming the pump. We avoid issuing insulin delivery commands to keep our implementation simple.
Ensure that the user always stays safe. Delay in delivering insulin does not pose any immediate threat to the user. However, delay in cancelling a wrongly issued insulin delivery command (such as one with a high insulin dose) can potentially kill the user. Thus, in the interest of keeping users safe, we violate the principle of not issuing commands to the pump in two specific cases of cancelling insulin delivery.
Disregard non-critical functionality in the interest of security. Some pump drivers incorporate an automatic retry mechanism when they’re unable to communicate with the pump. Such retry mechanisms complicate validation, since it becomes difficult to distinguish between duplicate commands issued as part of a retry versus duplicate commands issued as the result of a bug or an attack. We, thus, disregard such functionality by treating all duplicate commands as bugs or attacks.
@@ 5.2 Design Principles
5.3 Reference monitor design and implementation
The BioKernel invokes methods exposed by the reference monitor to register commands to be sent to the pump before invoking the relevant methods in the pump driver. The BioKernel also registers two callbacks with the reference monitor, one to be invoked on successfully receiving a response from the pump and the other in case of error.
The pump driver then updates its internal states, encodes the command and its relevant parameters and sends it to the reference monitor. The reference monitor then checks that the pump driver has not added new commands or modified the parameters before encrypting and sending them to the pump. If the reference monitor detects added or modified commands, it transitions the pump to operate in a default safe state and rejects all further commands from the driver.
We employ a bounded delay of one second to detect dropped commands. Concretely, the reference monitor detects dropped commands if it does not receive anything from the pump driver within one second of registering a command from the BioKernel. It treats dropped commands the same way as added or modified commands by transitioning the pump to a default safe state and rejecting all further commands. We transition the user out of this safe state and start accepting commands only after the user manually configures the BioKernel to trust the driver again.
The reference monitor operates differently on detecting dropped commands that deal with cancelling insulin delivery. Users will seek to cancel insulin delivery if they accidentally triggered an insulin delivery command or incorrectly set its parameters. Since excess insulin injection is dangerous, the reference monitor would automatically send these commands to the pump before transitioning the pump to the default state and rejecting all further commands. There are two commands that deal with cancelling insulin delivery: cancelDelivery and setting the suspend option on setTempBasal. cancelDelivery seeks to cancel a previous insulin delivery command and (if the delivery is still ongoing) and the suspend option on setTempBasal seeks to suspend all insulin delivery until the user resumes delivery.
@@ 5.3 Reference monitor design and implementation
OmniBLE does not incorporate automatic retry in its implementation. However, OmniBLE occasionally sends additional commands as part of its operation. For example, when sending a setTempBasal command, if OmniBLE’s internal states indicate that delivery is still in progress as part of a previous setTempBasal command, it first issues a cancelDelivery command to cancel that delivery. To remove ambiguity in terms of whether the cancelDelivery command was issued as part of setTempBasal, we enforce OmniBLE to always issue the cancelDelivery command. We evaluate GlucOS in terms of driver security in Section 10.4.
Overall, we employ straight-forward checks to detect manipulations from buggy/malicious pump drivers. However, the reference monitor’s design is more nuanced to handle the full range on insulin commands, missing out on any of which could potentially kill the user.
@@ 6 End-to-end security: Biological invariant
A key novelty in our overall approach is our introduction of a biological invariant that provides end-to-end verification. End-to-end verification in this context means that we verify the BioKernel’s glucose calculations against observations of the user’s metabolic state. If there is a large mismatch between our calculations and what we observe, then it means that our calculations are wrong. These deviations can happen from drastic changes in the user’s glucose absorption, like during exceptionally vigorous exercise or if the pump insertion site hits a vein and delivers insulin directly into the blood stream. Regardless of the cause, the biological invariant detects these conditions and prevents harm to the user.
The biological invariant closes the gap between verified software and human metabolism by incorporating a direct check against the user’s physiological data into the formal verification process. The biological invariant provides a runtime mechanism to detect when GlucOS’s calculations have become unsound. Specifically, the biological invariant compares the BioKernel’s theoretical calculation of the user’s blood glucose level against the actual observed blood glucose readings. Theoretical glucose levels provide an estimate of how the system thinks a given insulin dose will influence glucose levels. GlucOS uses its estimate of the human’s insulin sensitivity when calculating these doses. A significant divergence between the theoretical and observed values indicate a potential catastrophic failure mode, where GlucOS’s estimates of the user’s insulin sensitivity and calculations have diverged from the user’s real physiological state. The biological invariant allows the system to safely transition into a default safe state (Section 7). If the divergence between theoretical and observed glucose exceeds a threshold, indicating a loss of confidence in the model’s calculations, GlucOS discontinues automated insulin delivery and enters the safe state.
The biological invariant primarily focuses on resolving the undesirable situation where GlucOS believes that the person is less sensitive to insulin than they really are. In this case, the delivered insulin will remove more glucose than required, leading to potentially fatal low glucose levels. Consider the example of microbolusing when the user is above recommended glucose levels. If the system assumes a state where it believes the user is less sensitive to insulin than they really are and the user is 2x more sensitive, it will inject 2U of insulin when only 1U is required. While the system believes this will restore the user to safe levels, the actual glucose suppression will now push the user towards dangerously low glucose levels. Due to the higher clinical risk, we emphasize triggering the invariant in the undercalculation of insulin sensitivity case to transition to the default safe state. We set a threshold of 30 mg/dL for the maximum allowable divergence as shown in Figure 6 based on clinical evaluations showing detectability of significant glycemic deviations above this level[37, 32].
Figure 6: This figure shows an example of diverging differences between theoretical glucose and the observed glucose levels for a user. GlucOS will lose confidence in its calculationsand tranisition to a default safe state once it goes beyond a fixed difference threshold, where the biological invariant fails.
We formally model and verify the biological invariant using the Dafny language. The core aspects we specify include:
(a) A ghost variable representing the ground-truth physiological blood glucose level of the user. This models the actual, unobserved glucose state inside the human body that the system is trying to estimate and control.
(b) The theoretical calculation of the blood glucose level according to the system’s model. This captures the glucose prediction made by the system based on its assumptions about insulin dynamics, metabolism etc.
(c) The observed value of the blood glucose level from glucose readings from the user.
(d) The invariant property itself, which asserts that the absolute difference between the theoretical model glucose and the ground-truth physiological glucose cannot exceed a threshold of 30 mg/dL.
Our primary goal of verification in this segment is to make sure that there are no execution sequences that can violate the invariant without reaching a state where automated insulin delivery is disabled. This verification also ensures the biological invariant correctly captures our intention - if the system’s theoretical model deviates too far from the real physiological state, automated dosing is stopped to prevent potential over/under delivery of insulin. We also verify framing properties, showing the invariant’s state updates are precise and do not inadvertently modify other system variables. The verified Dafny specifications are then translated to Swift to integrate the biological invariant into the runtime system implementation for execution, with high assurance from the Dafny verification that they correctly implement the safe-guarding behavior in correspondence with the proven formal model.
Incorporating the biological invariant enables fundamental safety assurances not possible with traditional software verification alone. First, it constrains the set of possible erroneous system behaviors by restricting manual divergence between the theoretical model and actual physiological state. Section 10.5 shows that the biological invariant improves the system’s time-in-range glycemic control significantly compared to baseline when the system’s model and knowledge of user parameters becomes unsound. Second, it provides higher confidence of overall system reliability by directly incorporating physiological cross-checks into the verification process. We consider the biological invariant an important conceptual contribution that can extend the formal verification process to dynamically check and respond to deviations at runtime forming the complete bridge between the software and human physiology.
@@
Traditional software verification techniques like model checking [35] and deductive verification [24] have been explored in the past for hybrid system verification [23] that interact with the human body, but they do not completely account for physiological variability between individuals or model inaccuracies during the insulin delivery process.
7 Runtime failure management
GlucOS’s state machine consists of three operating states: (1) ML with safety operating in closed loop, (2) reactive safe model operating in closed loop and (3) manual dosing in closed loop. ML provides the tightest level of control in managing glucose but poses the largest immediate threat from excess insulin since it preemptively injects insulin by anticipating rise in glucose. The reactive safe model provides slightly lesser control, but poses milder immediate threat since it only injects insulin to account for measured excess glucose. Manual dosing offers the least amount of control but poses no threat since it only injects the minimum insulin needed for sustenance.
GlucOS predominantly operates in the state running ML (with safety) in closed loop. Certain real-world failures (such as missing CGM readings) transition GlucOS to operate in the state running the reactive safe model in closed loop. However, such failures are simple and we discuss ways to handle them in Appendix C. In this section, we focus on transitions pertaining to the manual dosing state.

### v2

@@ 3.2 Model principles
In contrast, predictive models have no restrictions since their outputs are constrained by the reactive safe model and our security mechanism. The ideal predictive model anticipates future metabolic states and adjusts insulin proactively to avoid glycemic imbalance.
In our evaluation, our predictive model outperforms the reactive safe model both in simulation (Section 8) and for the individual that uses our system for their real-world insulin dosing (Section 9), showing why it is important for automated insulin delivery systems to have a practical way to adopt ML.
@@ 4 Driver security: Withstanding malicious pump drivers
Pump drivers provide abstractions for insulin delivery applications to communicate with insulin pumps. Much like traditional device drivers, they have complex implementations since they help accomplish a wide range of tasks including pump setup/teardown, managing pump UX, handling insulin delivery/suspension, managing pump errors, accurately tracking insulin doses, and so on. Different pump drivers use different implementations with varying amounts of customizations and features to carry out these tasks.
While CVEs do not currently exist for pump drivers, there is strong possibility for bugs or vulnerabilities in their implementation given their complexity and the fact that they not been formally verified. Since drivers control insulin pumps, these vulnerabilities can harm people using them, even lead to the death. For example, consider a scenario where the user decides to inject 3 units (3U) of insulin using the BioKernel. The BioKernel would relay this as a command to the pump through abstractions provided by the pump driver. If the pump driver inadvertently or intentionally issues this command twice, 6U of insulin would be delivered to the user, which would put the user in danger by dosing too much insulin. We need a mechanism to verify that the pump driver is executing correctly.
Our security goals for GlucOS’s driver security are to protect users from inadvertent bugs in pump drivers and safeguard against potentially malicious pump drivers. To achieve these goals, we (1) introduce driver security logic in the BioKernel that interposes on pump driver-pump communications, (2) validate all commands against known high-level pump commands, and (3) detect and handle anomalies by stopping any added/modified commands and proactively handling dropping critical commands.
This approach allows people to choose between various pump drivers while maintaining security, balancing user flexibility and safety with automated insulin delivery. It also keeps our trusted computing base small, simplifying formal verification.
4.1 Driver security logic design
@@
Figure 4 provides an overview of our driver security logic. The BioKernel registers high-level commands, like “deliver 2U of insulin” with our driver security logic, and sends them to the pump driver. The pump driver converts these commands into I/O messages to program the pump. However, these I/O messages come back to the driver security logic, which checks to make sure that they are consistent with the previously registered high-level pump command, before forwarding them to the pump hardware.
We only validate commands for insulin delivery since they pose the biggest threats to the user. We do not validate communication pertaining to pump setup / UX since users can directly detect the respective manipulations. Figure 5 lists the different types of messages pertaining to insulin delivery exchanged between a pump driver and the OmniBLE insulin pump, which we use in our current implementation. We list other message types in Appendix B.
By interposing on the pump I/O message path, the driver security logic can ensure that the driver has not added, modified, or dropped any pump commands. Detecting added or modified commands is relatively straightforward. Our driver security logic in the BioKernel knows all high-level pump commands that it sends to the driver. Thus, we have a simple logic to confirm the corresponding I/O messages from the driver. The driver security logic stops any anomalous I/O messages before sending them to the pump.
Detecting dropped commands is more nuanced when the command that the driver drops is a command to cancel an ongoing insulin delivery. Users can cancel insulin delivery if they accidentally triggered an insulin delivery command or incorrectly set its parameters. Since excess insulin delivery is dangerous, the driver security logic automatically sends these cancel commands to the pump when it detects a dropped insulin cancellation command. There are two commands that deal with canceling insulin delivery: cancelDelivery and setting the suspend option on setTempBasal. The cancelDelivery command cancels a previous insulin delivery command (if the delivery is still ongoing) and the suspend option on the setTempBasal command suspends all insulin delivery until the user resumes delivery.
The most interesting lesson learned from our experience is in coping with dropped pump commands. In our original design, our goal was to ensure that our driver security logic only understood I/O commands and prevented any malicious I/O commands from going through. However, our driver security logic ended up needing to issue its own I/O commands to deal with dropped insulin delivery cancellation commands. This more proactive approach to driver security is different than previous approaches that also use specifications for runtime driver verification [69, 63, 44], which simply infer I/O semantics and do not issue their own I/O commands. In general, we avoid programming the pump directly from the driver security logic. But in this case, we decided to program the pump directly from the driver security logic for human safety.
When we detect a malicious driver, we transition the system into a safe manual mode where the individual needs to take over control of their own insulin delivery. They can still use the BioKernel to manage their insulin, but they would be unable to use the automated closed-loop algorithms. In our evaluation, our system had no false positives and detected all attacks (Section 8.2). See Appendix C for more details on how we maintain safety even when detecting an attack.
4.2 Driver security logic implementation
@@
OmniBLE does not incorporate automatic retry in its implementation. However, OmniBLE occasionally sends additional commands as part of its operation. For example, when sending a setTempBasal command, if OmniBLE’s internal states indicate that delivery is still in progress as part of a previous setTempBasal command, it first issues a cancelDelivery command to cancel that delivery. To remove ambiguity in terms of whether the cancelDelivery command was issued as part of setTempBasal, we enforce OmniBLE to always issue the cancelDelivery command.
@@ 5 End-to-end security: Biological invariant
A key novelty in our overall approach is our introduction of an end-to-end check that ensures the BioKernel’s glucose calculations against observations of the user’s metabolic state. This raises an interesting challenge for formal verification as it requires modeling the true biological state of the user – a parameter that is both unknown at runtime and distinct from measured parameters which are part of the software implementation. We refer to the true safety invariant we enforce on the user’s state as the biological invariant, and proving it is both a key part of our implementation and one of the key conclusions of our formal verification efforts.
@@
If there is a large mismatch between our calculations and what we observe, then it means that our calculations are wrong. These deviations can happen from drastic changes in the user’s glucose absorption, like during exercise when the body uses glucose without needing insulin [14], or if the pump insertion site hits a vein and delivers insulin directly into the blood stream [62]. Regardless of the cause, the end-to-end check detects these conditions and prevents harm to the user. For the purposes of the following example, suppose the end-to-end check is that the measured glucose levels are always within 30mg/dl of the predicted level; we refer to this as the implementation invariant that is enforced in an end-to-end way across the system:
$|g_{\mathsf{measured}}-g_{\mathsf{predicted}}|<30\mathsf{mg/dl}\quad(1)$
Naively, one might look at the $30\mathsf{mg/dl}$ on the right-hand side and think that the user’s true glucose levels are within $30\mathsf{mg}$ of the actual glucose levels. But this would be wrong! The problem is that the measured glucose level is subject to measurement error; the state-of-the-art Dexcom G7 CGM has a mean absolute relative difference of 8.2% from the true glucose level in the body [29]. For the sake of example, suppose that the measurement error on the CGM reading is around 5%; we refer to this as the CGM invariant:
$|g_{\mathsf{measured}}-g_{\mathsf{actual}}|<5\mathsf{mg/dl}\quad(2)$
and from this we can derive the true conclusion on the user’s glucose levels, which is what we refer to as the biological invariant. We stress that this invariant involves the parameter $g_{\mathsf{actual}}$ that is not available in software; it instead ensures a property (which we are able to formally prove) about the real level of glucose in the biological system. That statement we prove in the software is actually the following, not (1):
$|g_{\mathsf{measured}}-g_{\mathsf{predicted}}|<35\mathsf{mg/dl}\quad(3).$
When we run BioKernel on a real human, we don’t want to assume safety with respect to the measured glucose levels (invariant (1)); we actually want to ensure safety with respect to the real glucose levels in the human that is using the system (invariant (3)). Failing to distinguish these cases could be a severe software bug, and would jeopardize the end-to-end safety that we aim to provide by failing to account for CGM error. In the rest of the section, we describe the invariants that we enforce in the implementation (including invariant (1) as a simple case) and what additional invariants we prove in our formal verification development (including invariant (3)) below.
Implementation invariant.
In the implementation, we check that changes in glucose levels that we observe from CGM readings and the individual’s insulin delivery history are consistent with our theoretical calculations of how much the person’s glucose should change given these same inputs. If their glucose is dropping by more than our theoretical calculations show, then they are more sensitive to insulin then our system believes they should be. We set a threshold of 30 mg/dl per hour, for the maximum allowable divergence based on clinical evaluations showing detectability of significant glycemic deviations above this level [37, 34].
We take a three-tiered approach to dealing with deviations in our glucose change calculations. First, for small differences of less than 30 mg/dl, we accumulate errors and adjust the target glucose level to compensate, clamping the adjustment to 10 mg/dl. For example, if the individual’s insulin sensitivity is elevated and their target glucose level is 90 mg/dl, GlucOS will update their target to 100 mg/dl dynamically to adjust. Second, when the differences are 30 mg/dl or greater, we shut off insulin delivery. Our rationale is that the individual is more sensitive to insulin than our models believe they should be, thus we want to avoid overdosing. If their difference goes back to within the expected range, normal execution resumes. Third, if they remain at 30 mg/dl or more for more than two hours, we alert the individual and transition the system into a manual mode that requires the user to dose insulin manually.
Verification.
In the verification development (described in Section 7), we prove that the biological invariant holds, given error bounds set by the CGM manufacturer.
First, in the ported Dafny code, we locate the portions which enforce implementation invariants described above, including invariant (1). The code includes the two parameters $g_{\mathsf{measured}}$ and $g_{\mathsf{actual}}$, as these are available to the software implementation. Second, we model actual glucose by putting error bounds on our CGM readings, in the CGM invariant (2). This CGM invariant is inserted as a precondition on all functions that operate on CGM values. Third, we track the individual’s actual glucose level using a ghost variable in Dafny. Ghost variables are variables that are only available to the prover, these do not exist in the real software system. Fourth and lastly, we express as a postcondition the biological invariant (4), which asserts that the absolute difference between the theoretical model glucose and the ground-truth actual glucose cannot exceed a threshold of 35 mg/dl per hour. When the code is compiled, the Dafny verifier proves that the biological invariant holds on all inputs to the verified function.
Evaluation results.
Our evaluation shows that the biological invariant is effective at keeping humans safe when they face drastic changes in insulin sensitivity (Section 8.3), and that our thresholds of 30 mg/dl per hour and two hours were effective for the human that used GlucOS to manage their T1D without any false positives (Section 9.3).
Novelty.
@@
We beleive that having a parameter real biological state of the system is an important modeling decision; and it diverges, for example, from previous work which only models the biological data in simulation [71, 72]. Traditional software verification techniques like model checking [36] and deductive verification [27] have been explored in the past for hybrid system verification [26] that interact with the human body, but they do not completely account for physiological variability between individuals or model inaccuracies during the insulin delivery process. Also, researchers have used real-world invariants for security in the context of autonomous vehicles [18], but these were based on quad-copter physics not human physiology and did not include formal verification of their implementation.
6 Last line of defense: Humans
In this section, we describe our design and implementation for incorporating humans as a critical component of our multi-layered security system. While our algorithmic, system-level, and biological invariant defenses form the primary barriers against attacks, humans serve as the last line of defense, capable of mitigating attacks that manifest as glycemic imbalance.
When an individual’s glucose levels are out-of-range, they need to intervene to bring them back in range. Low glucose, or hypoglycemia, occurs from too much insulin, and there is no way to remove insulin that the automated insulin delivery system has already delivered. To recover from hypoglycemia, people need to eat sugar. Given that people with T1D are already active participants in managing glycemic imbalance, we focus on how we can use this fact to our advantage to create a more comprehensive security system.
Figure 6: GlucOS predicts hypoglycemia and notifies the individual before it happens so they can proactively prevent it.
We use predictive yet deterministic mobile notifications to alert people when they are likely to experience hypoglycemia before they reach it (Figure 6), as hypoglycemia presents the most immediate health risk and could be an indicator of an attack. Alerts for hypoglycemia are not new. All CGM software comes with non-predictive hypoglycemia alerts by default [2] and glucose prediction algorithms have been well covered in the literature [65, 70, 11]. Our contribution is using predictive alerts before someone reaches hypoglycemia as a security defense.
In a user study (Section 9.4), we reduced the amount of time six people spent in hypoglycemia as a result of using GlucOS, despite them already having CGM alerts set up prior to our study. These results demonstrate a reduction in hypoglycemia for the non-attack case, but the most dangerous successful attacks will result in hypoglycemia. Thus, these results suggest that predictive alerting can be an effective security measure.
Using a simple and deterministic algorithm is important since it runs as a part of our trusted computing base. Algorithmically, we predict hypoglycemia by doing a least squares fit for the last 20 minutes of CGM readings and extrapolate forward 15 minutes. We trigger alerts if our prediction shows hypoglycemia is likely in the next 15 minutes.
The most interesting finding from our study was that personalization was important. We provided people with the ability to set the prediction threshold and specify how often they wanted the alert to repeat while they remained out-of-range. Five of the six participants customized their alerts and the participant noted that the default values we picked coincided with their alerting preferences. In our first implementation we did not provide the ability to customize their security settings and we tried to make the usability vs security tradeoff algorithmically and automatically. But in the end, simplifying the system and allowing users to personalize their security policy was important for our participants.
Our finding has the potential to have broader implications for security in healthcare systems. Often, security systems favor uniformity and consistency [23, 28], but when managing people’s health we found that allowing people to balance alert fatigue and security increased engagement and potentially increased protection, based on each individual’s needs.